# Wildfire Risk Prediction: A Review

### v1

@@
Abstract
@@ 2.1.4 Fuel continuity
2.2 Climate and meteorology conditions
@@ 2.5 Wildfire historical records
MODIS data and products are also commonly used in wildfire risk prediction. For example, Bergado et al. (2021) used the ”Fire History Records of Fires primarily on Public Land” dataset provided by the State of Victoria, Department of Environment, Land, Water & Planning (DELWP) for their wildfire prediction research in Victoria, Australia. They believed that MODIS might overlook fires with smaller burn extents (Hantson et al., 2013). However, Abdollahi and Pradhan (2023) utilized both MODIS MCD64A1 and MOD14/MYD14 fire and thermal anomaly data, along with the NPWS Fire History - Wildfires and Prescribed Burns dataset provided by the NSW Department of Climate Change, Energy, the Environment, and Water, for their wildfire prediction study in the Gippsland region of Victoria, Australia. Ghorbanzadeh et al. (2019) obtained polygonal data for 34 wildfire areas through field surveys and evaluated them using MODIS data for the forestry area of Amol County in the Mazandaran province of northern Iran. Jaafari et al. (2019) and Nami et al. (2018) used historical data, field surveys, and validated them with MODIS hot spot products to enhance the reliability of historical data in their studies on the Zagros ecoregion in western Iran and the Hyrcanian ecological region in northern Iran, respectively. Kanwal et al. (2023) utilized the MODIS Fire Information for Resource Management System (FIRMS) as wildfire historical data for their research in Pakistan and set a confidence threshold of 60% to reduce false alarm errors. Zhang et al. (2021) used the monthly Global Fire Atlas (GFA) product from the Global Fire Emissions Database, covering the period 2003-2016, obtained from the GFA algorithm and 6 MCD64A1 burned area products with a 500 m resolution from MODIS. The GFA has a spatial resolution of 0.25° × 0.25°. They derived four seasonal fire maps, namely March to May (MAM), June to August (JJA), September to November (SON), and December to February (DJF), by averaging the monthly fire products. They analyzed wildfire susceptibility using a binary classification approach and constructed maps indicating the presence or absence of fires by reclassifying the four seasonal fire maps based on the 25th percentile.

### v2

@@
Wildfire Risk Prediction: A Review
Zhengsen Xu Jonathan Li Sibo Cheng Linlin Xu
Résumé
@@ 2.1.4 Fuel continuity
2.2 Climate and weahter conditions
@@ 2.5 Wildfire historical records
MODIS data and products are also commonly used in wildfire risk prediction. For example, Bergado et al. (2021) used the "Fire History Records of Fires primarily on Public Land" dataset provided by the State of Victoria, Department of Environment, Land, Water & Planning (DELWP) for their wildfire prediction research in Victoria, Australia. They believed that MODIS might overlook fires with smaller burn extents (Hantson et al., 2013). However, Abdollahi and Pradhan (2023) utilized both MODIS MCD64A1 and MOD14/MYD14 fire and thermal anomaly data, along with the NPWS Fire History - Wildfires and Prescribed Burns dataset provided by the NSW Department of Climate Change, Energy, the Environment, and Water, for their wildfire prediction study in the Gippsland region of Victoria, Australia. Ghorbanzadeh et al. (2019) obtained polygonal data for 34 wildfire areas through field surveys and evaluated them using MODIS data for the forestry area of Amol County in the Mazandaran province of northern Iran. Jaafari et al. (2019) and Nami et al. (2018) used historical data, field surveys, and validated them with MODIS hot spot products to enhance the reliability of historical data in their studies on the Zagros ecoregion in western Iran and the Hyrcanian ecological region in northern Iran, respectively. Kanwal et al. (2023) utilized the MODIS Fire Information for Resource Management System (FIRMS) as wildfire historical data for their research in Pakistan and set a confidence threshold of 60% to reduce false alarm errors. Zhang et al. (2021) used the monthly Global Fire Atlas (GFA) product from the Global Fire Emissions Database, covering the period 2003-2016, obtained from the GFA algorithm and 6 MCD64A1 burned area products with a 500 m resolution from MODIS. The GFA has a spatial resolution of 0.25° × 0.25°. They derived four seasonal fire maps, namely March to May (MAM), June to August (JJA), September to November (SON), and December to February (DJF), by averaging the monthly fire products. They analyzed wildfire susceptibility using a binary classification approach and constructed maps indicating the presence or absence of fires by reclassifying the four seasonal fire maps based on the 25th percentile.
@@ 5 Prediction Methods
5.1 Quantitative Statistics of Reviewed Papers
To provide a quantitative overview of commonly used methods, we analyzed 243 papers to determine the frequency of methods and features used. Additionally, we examined the spatial distributions of this research to highlight the fairness in research coverage.
5.1.1 Frequency of Different Methods
In this analysis, we categorized 247 models into statistical-based (S-based), traditional machine learning-based (TML-based), deep learning-based (DL-based), ensemble (E-based), experience-informed (E-informed), physics-informed (P-informed), and other methods, as illustrated in Fig. 1. Overall, the most frequently used method is traditional machine learning, appearing 115 times in our analysis. This is followed by deep learning methods and statistical-based methods, which were used 41 and 38 times, respectively.
Figure 1: The frequency of different methods in wildfire risk and spread prediction.
Within the traditional machine learning-based methods, random forest is the most preferred method, used 31 times in our analysis. This preference may stem from two aspects. Firstly, 31.25% of comparative studies indicate that random forest achieved state-of-the-art (SOTA) performance compared to other methods. Additionally, random forest offers explainability, which helps in identifying the importance of different features and aids decision-makers in understanding and intervening in factors influencing wildfire risk. Notably, most comparative studies did not include deep learning methods. Other traditional machine learning methods such as boosting, MLP, and SVM-based methods achieved SOTA performance 5, 2, and 1 times, respectively. Other methods, including decision tree, ANN, and Bayesian methods, were used approximately 11 times.
Statistical-based methods are also widely used due to their simplicity and ease of explanation. Among these methods, logistic regression is the most popular, utilized 20 times. GLM and WOE were each used 4 times. Additionally, 10 other regression-based methods were explored in various wildfire prediction studies.
Compared to shallow models, deep learning methods, particularly CNN and RNN, are also commonly used due to their high accuracy resulting from automatic deep feature extraction, despite their disadvantage in explainability. In our analysis, these two models were used 17 and 14 times, respectively, especially in combinations such as CNN and LSTM. Graph-based and vision transformer-based methods were used 4 and 1 times, respectively, to extract long-range dependencies given the complex fuel and environment of wildfire-prone areas. It is noteworthy that autoencoder-based methods were used 3 times, likely due to the lack of large-scale annotated wildfire data for deep learning model training.
Another approach to predicting wildfires is to combine various models and knowledge to enhance simulation, improve model performance, or reduce the data dependency of machine learning methods. In this section, we divided the remaining methods into four groups: combinations of different models, knowledge-informed methods, physics-informed models, and other models.
Ensemble methods are commonly used to predict wildfires. This research often involves combining several machine learning methods either to achieve final predictions or to form a complex wildfire risk prediction system. One common approach is to weight the predictions from different independent models to enhance accuracy (Ghorbanzadeh and Blaschke, 2018). For example, Rezaie et al. (2023) combined the group method of data handling (GMDH) and convolutional neural network (CNN) algorithms with biogeography-based optimization (BBO) and ant colony optimization (ACO) algorithms to improve predictive performance. Another approach involves building complex wildfire risk prediction systems. For instance, Ghorbanzadeh et al. (2019) combined machine learning (ML) methods with geographic information system multi-criteria decision-making (GIS-MCDM) methods to establish a forest fire susceptibility index and a social/infrastructure vulnerability index.
Compared to fully data-driven machine learning methods, 18 methods incorporated knowledge, such as observation bias, inductive bias, and learning bias (Karniadakis et al., 2021), to enhance model performance or reduce the reliance on large training data sets. For example, Troumbis et al. (2023) combined the theory of wildfire self-organized criticality (SOC) with modern statistics to provide a realistic explanation of the phenomenon’s evolution over recent decades and predict trends in Greece. Similarly, Bayani et al. (2023) used three-dimensional nonlinear vibration equations of power lines to generate a data set considering physical, structural, and meteorological parameters to address the problem of wildfires caused by power lines in California, USA. Machine learning techniques were then used to develop an agent model to quantify the risk of a single power line igniting a wildfire under extreme weather conditions. Additionally, Prestemon et al. (2002) used a fire risk function rather than a data-driven model to explore the relationship between wildfires and vegetation management, human land use, and climate factors in Florida counties.
Physics-informed models are commonly used in wildfire risk prediction tasks. Unlike knowledge-informed methods, these models are based on physical estimations rather than incorporating physical knowledge into other basic models. These methods are often used to directly estimate wildfire risk or simulate wildfire spread. For example, Liu et al. (2010) used a regional climate model to predict conditions in the northwestern United States. This model estimated precipitation, humidity, temperature, wind, soil, and vegetation conditions, which were then used to evaluate wildfire potential. Similarly, Mitsopoulos et al. used the minimum travel time algorithm to simulate wildfire risks under three different weather and fuel moisture conditions to better understand the interactions between wildfires and environmental and fuel conditions in Greece. Additionally, Gonzalez-Olabarria et al. (2023) employed an improved version of the Cell2Fire simulator, Cell2Fire_SB, to simulate the generation and spread of crown fires based on fuel models from Europe and the United States. Moreover, Lattimer et al. (2020) introduced machine learning into computational fluid dynamics (CFD)-based wildfire simulation, showing that prediction efficiency was 2-3 orders of magnitude faster than CFD simulations.
5.1.2 Research area distribution
Fig. 2 illustrates the geographic distribution and quantity of wildfire prediction research globally. The figure indicates that North America and Asia are the most concentrated regions for wildfire prediction studies. The United States has the highest number of studies, totaling 71, followed by China and Australia, each with 18 studies. Additionally, Canada, Iran, and India exhibit considerable research activity, with 11, 14, and 6 studies, respectively.
Figure 2: The spatial distribution of included research.
In Europe, the research is primarily concentrated in Western and Southern Europe. Spain, France, Greece, and Portugal have a notable number of studies, with 8, 5, 7, and 9 studies, respectively. Other countries such as Italy, Sweden, Serbia, and Turkey also have a smaller number of studies. Research on the African continent is relatively sparse, with South Africa, Algeria, Morocco, and Senegal each having one study. Similarly, South America has a limited number of studies, with Brazil and Chile each contributing one study. In Oceania, aside from Australia, no other countries have been recorded as having studies.
Furthermore, there are several studies categorized as global, totaling 18, and studies focused on the Mediterranean basin, totaling 6. These global and regional studies highlight the complexity and diversity of wildfire prediction across different climate zones and ecosystems.
Given the geographic distribution and quantity of research, future studies could focus on expanding research regions, promoting cross-regional studies, and enhancing data sharing and model validation. Specifically, although North America and Asia have concentrated research efforts, other regions such as Africa, South America, and parts of Europe have significantly fewer studies. Future research should aim to strengthen wildfire prediction studies in these underrepresented regions to address the geographic distribution gaps. Additionally, global and Mediterranean basin studies suggest that cross-regional research is crucial for understanding wildfire behavior under different ecosystems and climatic conditions. Enhancing international collaboration to conduct more comprehensive cross-regional studies can provide more robust prediction models and methods. Finally, many current studies rely on region-specific datasets, limiting the generalizability and applicability of the models. Future efforts should focus on promoting data sharing and openness, establishing unified data platforms, and conducting model validation across different regions to improve the reliability and applicability of prediction models.